OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. (x - 5) 2 + (y + 3) 2 = 25 center = ( , ) radius =

OpenStudy (2bornot2b):

center = ( 5 , -3 ) radius = 5

OpenStudy (ash2326):

standard equation of a circle = (x-a)^2+ (y-b)^2= r^2 it's center is (a, b) and radius= r we have the equation (x - 5)^2 + (y + 3)^2 = 25 so center= (5,-3) radius= 5